# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v3

@@ 6 State
Following Fig. 6, the pose estimation method primarily uses the generative model to generate a Normalized Object Coordinate Space (NOCS) map (Cai et al. (2024)). The scene reconstruction method uses features from the VLM image encoder, such as CLIP (Gao et al. (2024)). These features serve as supervision labels for NeRF. This approach ensures that the reconstructed scene contains semantic information (Kerr et al. (2023)). The VLM image encoder can extract features from multi-view images. These features can then be lifted to 3D space (Hong et al. (2023b)).
@@ 7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (Trivedi et al. (2021). However, a robot capable of generating code for multiple tasks, rich knowledge across various domains is essential (Ellis et al. (2023)). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (Chen et al. (2021); Austin et al. (2021)). Code-As-Policy (Liang et al. (2023)) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capability, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. In PROGPROMPT (Singh et al. (2023)), assertions are added to the generated code. When executing assertions, environment state feedback is obtained to check if the environment satisfies the pre- and post-conditions of the task.
@@ 7.2 VLATP
Utilizing foundation models trained on existing large-scale internet datasets enables the direct perception of observation images and outputting target poses. Instruct2Act (Huang et al. (2023b)) utilizes CLIP and SAM to identify manipulated objects within observation images and outputs the 3D position of these manipulated objects from 2D image. DALL-E-Bot (Kapelyukh et al. (2023)) employs DALL-E to generate target images for tasks and generates target poses for manipulation by combining the target image with the observation image. ReKep (Huang et al. (2024)) uses VFM and VLM to extract relational keypoint constraints from language instructions and RGBD observations. It then applies an optimization solver to generate a series of end-effector poses.
@@
RFMs based on the above work are also under development. LEO (Huang et al. (2023a)) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023b)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. Xu et al. (2024) considers the task goal, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the target pose. However, the target pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point.
@@ 7.3 VLADP
RoboNet (Dasari et al. (2019)) establishes an initial pool of 15 million video frames from 7 different robot platforms, aiming to learn a generalizable model for vision-based robotic manipulation through an end-to-end approach. This learning method closely mimics human learning, eliminating the need for calibration and 2D-to-3D mapping to understand spatial 3D coordinates. Due to imagination, RoboNet generates predicted 2D images of the next movement and estimates which pixel the robot should move to in the next step, ultimately converting it into the action the robot should take. This concept of imagination is applied in both 3D-VLA (Zhen et al. (2024)) and SuSIE (Black et al. (2023)). However, RobotNet can only achieve one-model corresponding to one-task. One-model corresponding to one-task methods have also achieved significant progress. ACT (Zhao et al. (2023a)) uses action chunks to reduce compounding errors. Diffusion policy (Chi et al. (2023)) applies the idea of diffusion to visuo-motor control, tackling challenges such as action multimodality and sequential correlation to handle high-dimensional action sequences.
@@
Utilizing pre-trained VLMs (Zhang et al. (2024a)) for fine-tuning to construct RFMs is considered efficient. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating delta pose as tokens. However, this approach necessitates a lot of data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). However, the current Open X-Embodiment dataset faces the heterogeneity dataset challenge. Octo (Team et al. (2024)), RDT-1B (Liu et al. (2024d)), HPT (Wang et al. (2024c)) and PI (Black et al. (2024)) propose multi-module networks to address this issue. RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses and human intervention can enable robots to adjust trajectories based on human language instruction. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023c)) and OpenVLA (Kim et al. (2024)) emerge, effectively enabling a robot manipulation policy with VLMs.
@@ 2 NVIDIA
7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, effectively exploring, reusing experience data, and designing reward functions (Kober et al. (2013)). Foundation models have demonstrated the emergence of common sense reasoning, the ability to propose and sequence sub-goals, and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capability of foundation models to address the challenges faced by reinforcement learning. RobotGPT (Jin et al. (2024)) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, many studies explore the use of foundation models to solve challenges like long-horizon problems and effectively exploring and designing reward functions.
Summary
@@ 8 Manipulation Data Generation
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (Wang et al. (2023b)). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. Matas et al. (2018) trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge. Regardless of whether it’s in a real or simulated environment, improving the efficiency of the existing dataset is essential. The mainstream approach is dataset augmentation.
@@ 8.2 Scene and Demonstration Generation
Simulation scenes can be created manually. However, this approach is time-consuming and labor-intensive. As a result, automated or semi-automated scene generation methods are more preferred (Deitke et al. (2022)). Two methods can be used. Real-to-Sim method converts real scenes to simulation. Automated generation method generates simulation scenes automatically without real-world observation. Real-to-Sim method can accurately mimic the real world, but it limits the diversity of scenes. The automated generation method can create more diverse scenes and increase the variety of collected demonstrations.
@@
The Real-to-Sim method directly refers to a digital twin. The Real-to-Sim method utilizes 3D reconstruction technology (in Sec. 6) or inverse graphics (Chen et al. (2024c)) to create the real-world scene in a virtual environment (Torne et al. (2024)). But, 3D reconstruct scene is static environment where objects lack real-world physical properties, such as material, mass and friction coefficients, and are non-interactive (Torne et al. (2024)). The inverse graphics method, such as URDFormer (Chen et al. (2024c)), directly generates interactive simulation environment and articulated objects from input RGB image. Compared to 3D-generated methods, it reduces human involvement and produces interactive simulation environment. However, it lacks physical plausibility and fails to address the mismatch between the generated object’s physical properties and the real world. Previous research on physical parameter identification of real-world objects can be categorized into estimation from interaction (Seker and Kroemer (2024); Bohg et al. (2017); Xu et al. (2019)), estimation from demonstrations (Torne et al. (2024)). PHYSOBJECTS (Gao et al. (2023)) fine-tunes a VLM to estimate physical object attributes, such as material, from observation.
@@
GRS (Zook et al. (2024)) employs SAM2 for object segmentation from RGBD image and utilizes VLMs to describe and match objects with simulation-ready assets. This approach combines the strengths of 3D reconstruction and inverse graphics methods. It ensures the credibility of 3D-generated methods and allowing objects in the scene to interact. However, it is impossible for the assert dataset to fully cover objects in the real world. ACDC (Dai et al. (2024)) defines a digital cousin concept. Unlike a digital twin, it does not directly replicate a real-world counterpart. However, it retains similar geometric and semantic features by using similar asserts when the assert dataset does not include real-world objects. Constructing an interact assert dataset often requires manual design by the creator or human-assisted interactive object generation.
@@
The automated generation methods can be categorized into rule-based and learning-based approaches. For instance, ProcTHOR (Deitke et al. (2022)) introduces a procedural generation pipeline for interactive scenes using rule-based constraints and statistical priors. However, the generated scenes often rely on pre-defined priors, resulting in unrealistic outcomes that hinder agent learning (Khanna et al. (2024)). To address this, PHYSCENE (Yang et al. (2024)) incorporates physical collision avoidance, object layouts, interactivity, and reachability metrics into a diffusion model. This approach enhances the physical plausibility and interactivity of generated scenes. Due to the prior knowledge of foundation models, there are current efforts to use foundation models for scene construction. RoboGen (Wang et al. (2023d)) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (Wang et al. (2023b)) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library. However, using foundation models to automate the generation of scene’s physical plausibility still relies on VLM for judgment.
@@
To achieve scene diversity, RoboCasa (Nasiriany et al. (2024)) uses Madjourney to generate texture images for kitchen scenes. At the same time, many studies focus on the automatic generation of assets, such as zero-1-to-3 (Wang et al. (2023d)), Luma.ai (Nasiriany et al. (2024)), LLaMA-Mesh (Wang et al. (2024d)). The modeling of the interaction environment above primarily focuses on articulated object modeling. Articulated objects can be created manually by designers or generated using procedural (Jiang et al. (2022); Liu et al. (2023b); Zhang et al. (2023d)) or human-assisted interactive methods (Torne et al. (2024)) after 3D scanning. They can also be generated automatically through inverse graphics (Chen et al. (2024c)). However, current automated methods for generating articulated object assets are limited to objects with few rotational joints. Real2Code (Mandi et al. (2024)) fine-tunes a CodeLlama model to process visual observation descriptions and then outputs joint predictions. This enables Real2Code to reconstruct complex articulated objects with up to 10 parts.
@@
To collect demonstrations in simulations, different approaches can be used based on task complexity. For simple tasks, like a two-finger gripper picking up a cube, a hard-coding method (Wang et al. (2022b)) can be used. However, for more complex tasks, remote teleoperation (Chen et al. (2024a)) or skill library (Ha et al. (2023)) should be employed. Building skill library can be done using reinforcement learning or gradient optimization methods. RoboGen (Wang et al. (2023d)) shows that gradient-based trajectory optimization is better for fine-grained manipulation tasks with soft bodies, like shaping dough into a specific form. On the other hand, reinforcement learning and evolutionary strategies are more effective for contact-rich tasks and continuous interactions with other components in the scene.

### v4

@@ 6.2 Pose Estimation
Following Fig. 6, VLM and VFM assist implicit 3D reconstruction by generating relevancy maps that include semantic information (Kerr et al. (2023)). They can also be employed in 2D-to-3D lifting to extract 3D features, encompassing texture, semantic, and spatial information (Hong et al. (2023b)). VGM aids in generating edited images and modifying 3D scenes based on these edited images (Haque et al. (2023)). FreeZe (Caraffa et al. (2024)) achieves state-of-the-art result in pose estimation by extracting discriminative features through 2D-to-3D lifting and LMM. However, the performance is limited by the foundation models themselves. For example, CLIP is constrained by the bag-of-words limitation in semantic scene understanding (Kerr et al. (2023)).
Figure 6: Foundation Models for State. The foundation models have three main applications in 3D reconstruction: 3D open-set semantic scene understanding, lifting 2D features to 3D space, and implicit 3D editing. In 3D open-set semantic scene understanding, the main pipeline is to use image features extracted by the VFM encoder and VLM image encoder as input for NeRF. Then, semantic text features extracted by the VLM language encoder are used in conjunction with the image features from NeRF to generate a relevancy map through a relevancy extractor (Kerr et al. (2023)). This relevancy map can support downstream tasks, such as semantic scene completion and object localization (Ha and Song (2022)). As for lifting, using the VLM image encoder to extract features from 3D data multi-view images and lift them into 3D features can incorporate semantic information into the 3D features. The lifting methods include direct reconstruction, gradSLAM, and Neural Field (Hong et al. (2023b)). For implicit 3D editing, the current mainstream pipeline is to input the image rendered by NeRF and the editing prompt into the VGM to generate the updated image. The updated image is then fed back into NeRF for training, modifying NeRF’s radiance field representation of the 3D scene (Haque et al. (2023)). Pose estimation with foundation models achieves state-of-the-art results (Caraffa et al. (2024)). The main method is 2D-lifting-3D. It extracts texture features from the object model and observation RGBD image. LMM extracts geometric features from the object model and observation RGBD image. The fused features are then used to estimate the 6D pose through 3D-3D correspondences.
@@ 7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (Trivedi et al. (2021). However, a robot capable of generating code for multiple tasks, rich knowledge across various domains is essential (Ellis et al. (2023)). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (Chen et al. (2021b); Austin et al. (2021)). Code-As-Policy (Liang et al. (2023)) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capability, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. In PROGPROMPT (Singh et al. (2023)), assertions are added to the generated code. When executing assertions, environment state feedback is obtained to check if the environment satisfies the pre- and post-conditions of the task.
@@ 7.2 VLATP
Utilizing foundation models trained on existing large-scale internet datasets enables the direct perception of observation images and outputting target poses. Instruct2Act (Huang et al. (2023b)) utilizes CLIP and SAM to identify manipulated objects within observation images and outputs the 3D position of these manipulated objects from 2D image. DALL-E-Bot (Kapelyukh et al. (2023)) employs DALL-E to generate target images for tasks and generates target poses for manipulation by combining the target image with the observation image. ReKep (Huang et al. (2024a)) uses VFM and VLM to extract relational keypoint constraints from language instructions and RGBD observations. It then applies an optimization solver to generate a series of end-effector poses.
@@
RFMs based on the above work are also under development. LEO (Huang et al. (2023a)) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023c)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. Xu et al. (2024) considers the task goal, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the target pose. However, the target pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point.
@@ 7.3 VLADP
RoboNet (Dasari et al. (2019)) establishes an initial pool of 15 million video frames from 7 different robot platforms, aiming to learn a generalizable model for vision-based robotic manipulation through an end-to-end approach. This learning method closely mimics human learning, eliminating the need for calibration and 2D-to-3D mapping to understand spatial 3D coordinates. Due to imagination, RoboNet generates predicted 2D images of the next movement and estimates which pixel the robot should move to in the next step, ultimately converting it into the action the robot should take. This concept of imagination is applied in both 3D-VLA (Zhen et al. (2024)) and SuSIE (Black et al. (2023)).
RobotNet can only achieve one-model corresponding to one-task. One-model corresponding to one-task methods have also achieved significant progress. ACT (Zhao et al. (2023a)) uses action chunks to reduce compounding errors. Diffusion policy (Chi et al. (2023)) applies the idea of diffusion to visuo-motor control, tackling challenges such as action multimodality and sequential correlation to handle high-dimensional action sequences.
@@
Utilizing pre-trained VLMs (Zhang et al. (2024b)) for fine-tuning to construct RFMs is considered efficient. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating delta pose as tokens. However, this approach necessitates a lot of data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). However, the current Open X-Embodiment dataset faces the heterogeneity dataset challenge. Octo (Team et al. (2024)), RDT-1B (Liu et al. (2024d)), HPT (Wang et al. (2024c)) and PI (Black et al. (2024)) propose multi-module networks to address this issue. RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses and human intervention can enable robots to adjust trajectories based on human language instruction. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023d)) and OpenVLA (Kim et al. (2024)) emerge, effectively enabling a robot manipulation policy with VLMs.
@@
7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, effectively exploring, reusing experience data, and designing reward functions (Kober et al. (2013)). Foundation models have demonstrated the emergence of common sense reasoning, the ability to sequence sub-goals and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capability of foundation models to address the challenges faced by reinforcement learning. RobotGPT (Jin et al. (2024)) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, many studies explore the use of foundation models to solve challenges like long-horizon problems and effectively exploring and designing reward functions.
Summary
@@ 8 Manipulation Data Generation
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (Wang et al. (2023c)). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. Matas et al. (2018) trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge. Regardless of whether it’s in a real or simulated environment, improving the efficiency of the existing dataset is essential. The mainstream approach is dataset augmentation.
@@ 8.2 Scene and Demonstration Generation
Simulation scenes can be created manually. However, this approach is time-consuming and labor-intensive. As a result, automated or semi-automated scene generation methods are more preferred (Deitke et al. (2022)). Two methods can be used. Real-to-Sim method converts real scenes to simulation. Automated generation method automatically generates simulation scenes without real-world observation. Real-to-Sim method can accurately mimic the real world, but it limits the diversity of scenes. The automated generation method can create more diverse scenes and increase the variety of collected demonstrations.
@@
The Real-to-Sim method directly refers to a digital twin. The Real-to-Sim method utilizes 3D-reconstruction technology (in Sec. 6) or inverse graphics (Chen et al. (2024c)) to create the real-world scene in a virtual environment (Torne et al. (2024)). But, 3D reconstruct scene is static environment where objects lack real-world physical properties, such as material, mass and friction coefficients, and are non-interactive (Torne et al. (2024)). The inverse graphics method, such as URDFormer (Chen et al. (2024c)), directly generates interactive simulation environment and articulated objects from input RGB image. Compared to 3D-reconstruction methods, it reduces human involvement and produces interactive simulation environment. However, it lacks physical plausibility and fails to address the mismatch between the generated object’s physical properties and the real world. Previous research on physical parameter identification of real-world objects can be categorized into estimation from interaction (Seker and Kroemer (2024); Bohg et al. (2017); Xu et al. (2019)), estimation from demonstrations (Torne et al. (2024)). PHYSOBJECTS (Gao et al. (2023)) fine-tunes a VLM to estimate physical object attributes, such as material, from observation.
@@
As for the application of foundation models in Real-to-Sim methods, GRS (Zook et al. (2024)) employs SAM2 for object segmentation from RGBD image and utilizes VLMs to describe and match objects with simulation-ready assets. This approach combines the strengths of 3D-reconstruction and inverse graphics methods. It ensures the credibility of 3D-reconstruction methods and allowing objects in the scene to interact. However, it is impossible for the assert dataset to fully cover objects in the real world. Constructing an interact assert dataset often requires manual design by the creator or human-assisted interactive object generation. ACDC (Dai et al. (2024)) defines a digital cousin concept. Unlike a digital twin, it does not directly replicate a real-world counterpart. However, it retains similar geometric and semantic features by using similar asserts when the assert dataset does not include real-world objects. As for object pose, depth cameras are commonly used, but they struggle to capture reflective surfaces accurately. This limits their use in the wild. To address this issue, ACDC uses Depth-Anything-v2 (Yang et al. (2024a)), a state-of-the-art monocular depth estimation model, to estimate the depth map.
@@
The automated generation methods can be categorized into rule-based and learning-based approaches. For instance, ProcTHOR (Deitke et al. (2022)) introduces a procedural generation pipeline for interactive scenes using rule-based constraints and statistical priors. However, the generated scenes often rely on pre-defined priors, resulting in unrealistic outcomes that hinder agent learning (Khanna et al. (2024)). To address this, PHYSCENE (Yang et al. (2024b)) incorporates physical collision avoidance, object layouts, interactivity, and reachability metrics into a diffusion model. This approach enhances the physical plausibility and interactivity of generated scenes.
Due to the prior knowledge of foundation models, there are current efforts to use foundation models for scene construction. RoboGen (Wang et al. (2023e)) utilizes LLM to generate relevant assets, asset sizes, asset configuration, scene configuration based on the task proposals and use text-to-image-to-3D generation to create the corresponding assets. These assets are imported into the simulator to generate the appropriate scene. Finally, using VLM for task-specific scene verification. GenSim (Wang et al. (2023c)) uses LLMs to generate new task and task scenario codes based on the pre-cached scene codes in a task library. However, using foundation models to automate the generation of scene’s physical plausibility still relies on VLM for judgment.
@@
The Real-to-Sim method and the Automated generation method both rely on 3D assets. The diversity of 3D assets determines the variety of scenes. RoboCasa (Nasiriany et al. (2024)) uses Madjourney to generate texture images for kitchen scenes. At the same time, many studies focus on the automatic generation of assets, such as zero-1-to-3 (Wang et al. (2023e)), Luma.ai (Nasiriany et al. (2024)), LLaMA-Mesh (Wang et al. (2024d)), Trellis (Xiang et al. (2024)). The modeling of the interaction environment above primarily focuses on articulated object modeling. Articulated objects can be created manually by designers or generated using procedural (Jiang et al. (2022); Liu et al. (2023b); Zhang et al. (2023e)) or human-assisted interactive methods (Torne et al. (2024)) after 3D scanning. They can also be generated automatically through inverse graphics (Chen et al. (2024c)). However, current automated methods for generating articulated object assets are limited to objects with few rotational joints. Real2Code (Mandi et al. (2024)) fine-tunes a CodeLlama model to process visual observation descriptions and then outputs joint predictions. This enables Real2Code to reconstruct complex articulated objects with up to 10 parts.
@@
To collect demonstrations in simulations, different approaches can be used based on task complexity. For simple tasks, like a two-finger gripper picking up a cube, a hard-coding method (Wang et al. (2022b)) can be used. However, for more complex tasks, remote teleoperation (Chen et al. (2024a)) or skill library (Ha et al. (2023)) should be employed. Building skill library can be done using reinforcement learning or gradient optimization methods. RoboGen (Wang et al. (2023e)) shows that gradient-based trajectory optimization is better for fine-grained manipulation tasks with soft bodies, like shaping dough into a specific form. On the other hand, reinforcement learning and evolutionary strategies are more effective for contact-rich tasks and continuous interactions with other components in the scene.